# Eternal Universes

### v1

@@ III Eternal inflating universe
Our first example of an eternal universe was introduced in Lesnefsky et al. (2023b), and analyzed in detail in Easson and Lesnefsky (2024):
@@
The cosmological evolution is depicted in Fig. 1. Shown are the scale factor $a(t)$, Hubble parameter $H$ (amplified by a factor of 100) and co-moving Hubble radius $1/aH$. When the co-moving Hubble radius is decreasing the spacetime is inflating–in this case it is decreasing for all $t$.
Figure 1: Evolutionary behavior for (3.11). The scale factor is plotted in red. The Hubble parameter (green). The co-moving Hubble radius (blue). Model parameters: $a_{0}=c=1/2$, $\alpha=1$.
@@
Figure 2: Curvature invariants computed from (3.11). The Ricci scalar is plotted in green and the Kretchmann scalar is plotted in red. Model parameters: $a_{0}=c=1/2$, $\alpha=1$.
@@
Figure 3: Energy conditions from (III). Plot of energy density $\rho$ (red), $\rho+p$ (yellow), $|p|$ (blue) and $\rho+3p$ (green). Model parameters: $a_{0}=c=1/2$, $\alpha=1$.
All of the energy conditions are violated. NEC violation is identified by the dipping of the yellow curve ($\rho+p$) below the vertical axis, both here, and in all of the proceeding energy condition plots. The NEC asymptotically approaches saturation in the early and late universe since $\rho+p=-2\operatorname{sech}^{2}{t}\rightarrow 0$ as $t\rightarrow\pm\infty$. It is easy to see the ANEC (Eq. 1.1) is also violated in this scenario. We will see that the ANEC violation is not a generic feature of eternal models. In the late universe, the SEC remains violated allowing continued acceleration, as indicated by the negativity of the green curve.
@@ IV.1 Transcendental bounce
The above scale factor describes a universe that contracts, bounces and then expands (similar to the familiar case of closed de Sitter).  [^10] A solution with this form was also found in a non-local higher-derivative gravity in Biswas et al. (2006, 2012), which was argued to be free of pathologies, which if true would be quite remarkable given the level of energy condition violation (see Fig. 6). The Hubble parameter is
@@
$\int^{t}\frac{a\left(\zeta\right)d\zeta}{\sqrt{\left(a\left(t\right)\right)^{2}+1}}=\alpha\operatorname{arctanh}\left(\frac{a_{0}\sinh\left(\frac{t}{\alpha}\right)}{\sqrt{1+a_{0}^{2}+a_{0}^{2}\sinh\left(\frac{t}{\alpha}\right)^{2}}}\right)\,,$ (4.19)
@@
[…]
@@
Figure 4: Evolutionary behavior for Eq. 4.16. The scale factor (red). The Hubble parameter amplified by an order of magnitude (green). The co-moving Hubble radius amplified by an order of magnitude (blue). Parameters: $a_{0}=\alpha=1$.
@@
Figure 6: Energy conditions from Eq. IV.1. Plot of energy density $\rho$ (red), $\rho+p$ (yellow), $|p|$ (blue) and $\rho+3p$ (green). Model parameters: $a_{0}=\alpha=1$.
@@ IV.2 Polynomial bounce
Corollary 4.
In General Relativity, every smooth, non-constant scale factor $a(t)$ of a geodesically complete FRW spacetime must undergo accelerated expansion for at least some period of time.
@@
Building on O’Neill (1983) Prop 12.15, let $\mathcal{M}(k,a)=I\times_{a}S$ be a FRW spacetime with Lorentzian manifold $\mathcal{M}$, smooth scale factor $a$ on an open interval $I$ in $\mathbb{R}^{1}_{1}$ and connected three-dimensional Riemannian manifold $S$ of constant curvature $k\in\{-1,0,1\}$. If $H_{0}=\dot{a}(t_{0})/a(t_{0})>0$ for some $t_{0}$, and the SEC is satisfied, $\rho+3p>0$, then $I$ has an initial endpoint $t_{*}$ with $t_{0}-H_{0}^{-1}<t_{*}<t_{0}$, and either (1) $\dot{a}>0$, or (2) $a$ has a maximum point after $t_{0}$, and $I$ is a finite interval $(t_{*},t^{*})$.
This is easily seen from the Friedmann acceleration equation, $3\ddot{a}/a=-4\pi G(\rho+3p)$, which implies for matter obeying the SEC that $\ddot{a}<0$. Thus, except at $t_{0}$, the graph of $a$ lies below its tangent line at $t_{0}$. This line represents the linear approximation of the scale factor at $t_{0}$, expressed by: $A(t)=a(t_{0})+H_{0}\dot{a}(t_{0})(t-t_{0}).$ Given that $H_{0}>0$, as time $t$ regresses from $t_{0}$, the scale factor $a(t)>0$ must encounter a singularity at some $t_{*}$ before it can reach the zero of $A(t)$ at $t_{0}-H_{0}^{-1}$. Hence, assuming existence of a FRW spacetime that is geodesically complete (and non-trivially evolving), it must violate the SEC, which implies $\ddot{a}>0$, i.e. it must positively accelerate (inflate) for at least some amount of time. ∎
The above Corollary for generalized FRW metrics has no known proof, yet we offer:
Conjecture 5.
Every smooth, non-constant scale factor $a(t)$ of a geodesically complete GFRW spacetime must inflate for at least some period of time.
The conjecture is supported by the following discussion. Considering the sufficient condition that any scale factor $a\left(t\right)$ with strictly positive infimum is geodesically complete under Thm. 1, one can examine the implication this has upon the concavity - and hence possible inflationary behaviour - of $a\left(t\right)$. For the (assumed smooth and non-trivial) scale factor function to have strictly positive infimum it is difficult to have negative concavity, and hence never inflate, for all $\mathbb{R}$. Unless there is a change in concavity and the scale factor undergoes inflation, $a\left(t\right)$ will fail to be strictly positive and a singularity will occur.
Lending credence due to the lack of formal proof of the conjecture is the following construction: consider a function which is concave down at all times, but, as discussed in Appendix D, is not strictly positive over its domain, say $-t^{2}+1$. One can then utilize the fact that any two connected intervals of $\mathbb{R}$, including $\mathbb{R}$ itself are diffeomorphic, and compose a (smooth) diffeomorphism between $\mathbb{R}$ and an interval where $-t^{2}+1>0$, say $\left(-\nicefrac{{1}}{{4}},\nicefrac{{1}}{{4}}\right)$. Naively, one may believe that this is a counter-example to the aforementioned conjecture, however closer inspection of the diffeomorphism is required.
Although the above conjecture has no known formal proof, a particular class of scale factors does:
Corollary 6.
Let $\mathbb{R}^{1}_{1}\times_{p}\Sigma$ be a GFRW with scale factor $p\left(t\right)\in\mathbb{R}\left[t\right]$ be an “appropriate” polynomial of degree $2n$ with $n\geq 1$, then $\mathbb{R}^{1}_{1}\times_{p}\Sigma$ undergoes inflation.
@@
The scale factor $p\left(t\right)$ is assumed to be non-constant. Said polynomial has at most $12n-$ 1 critical points, thus any local inflection changes occur over a compact finite interval in $\mathbb{R}$. However, the left-most and right-most critical points cannot be a local maximum - and thus have downward inflection - because, by the assumptions inherent in “appropriate” polynomials, the tails of $p\left(t\right)$ increase without bound. Thus the tails of the scale factor must have upward inflection and hence the scale factor is inflationary. ∎
Considering Conjecture 5, Corollary 6 and Defn. 1 we find complete spacetimes must be inflationary Easson and Lesnefsky (2024). [^11] We pause here to acknowledge the celebrated work of Alexei Starobinsky who was motivated to find non-singular cosmological solutions by having an early-time accelerated expansion Starobinsky (1980), opening the window for the theory of the inflationary universe, and aligning with Cor.4.
IV.2.1 Quadratic model
@@
Figure 9: Energy conditions from Eq. IV.2.1. Plot of energy density $\rho$ (red), $\rho+p$ (yellow), $|p|$ (blue) and $\rho+3p$ (green). Model parameters: $a_{0}=b=1$, $c=2$.
@@
In this model, and others we consider, the NEC is violated during a finite time interval, $-t_{-}<t<t_{+}$. One may break up the ANEC integral into the asymptotic regions $t<-t_{-}$, $t>t_{+}$ where other energy conditions ( e.g. NEC, WEC and DEC) are satisfied, and the NEC violating region. In this case the ANEC condition Eq. 1.1 may be recast as Giovannini (2017):
[…]
@@
In this case the ANEC is found to be satisfied since,
[…]
In the late universe, only the SEC is violated allowing continued acceleration, as indicated by the negativity of the green curve. This too is an eternally inflating, bouncing cosmology similar to the behavior of Eq. 4.16, but it is better-behaved with respect to energy conditions as the NEC violation may be restricted to a moment in the very early universe  [^12] A modern example of such a bounce with momentary NEC violation is the S-brane model Brandenberger et al. (2014)., and the ANEC Eqs. 1.1 and 4.30 is satisfied.
IV.2.2 Models from limiting curvature
One method for building such non-singular bounces stems from the limiting curvature hypothesis (LCH) which stipulates the existence of minimal length (or maximal curvature) scale in nature, presumably somewhere near the Planck length Markov (1982, 1987), although one may introduce new physics at lower energy scales using this method. By construction, curvature invariants are limited in an action resulting in non-singular solutions. The LCH has been used to successfully construct many non-singular cosmological solutions Mukhanov and Brandenberger (1992); Brandenberger et al. (1993); Moessner and Trodden (1995); Brandenberger (1995); Brandenberger et al. (1998); Easson and Brandenberger (1999); Easson (2003a); Chamseddine and Mukhanov (2017a) in a variety of settings, as well as non-singular black hole solutions Frolov et al. (1989, 1990); Morgan (1991); Trodden et al. (1993); Easson (2003b, 2018); Chamseddine and Mukhanov (2017b); Brandenberger et al. (2021); Frolov and Zelnikov (2022).
@@ IV.2.2 Models from limiting curvature
for constants $c>0,b>0$, which is capable of smoothly connecting two matter dominated phases having $a\propto t^{2/3}$ via a NEC violating bounce at $t=0$ where $\ddot{a}>0$. The schematic behavior of the cosmological quantities and the energy conditions are very similar to the polynomial model analyzed above; however, in the present case, even the SEC is obeyed in both the early and late (non-accelerating) universe. The ANEC Eqs. 1.1 and 4.30 is satisfied:
[…]

### v2

@@ III Eternal inflating universe
Our first example of an eternal universe was introduced in Lesnefsky et al. (2023), and analyzed in detail in Easson and Lesnefsky (2024):
@@
The cosmological evolution is depicted in Fig. 1. Shown are the scale factor $a(t)$, Hubble parameter $H$ (amplified by an order of magnitude) and co-moving Hubble radius $1/aH$. When the co-moving Hubble radius is decreasing the spacetime is inflating–in this case it is decreasing for all $t$.
Figure 1: Evolutionary behavior for (3.11). The scale factor is plotted in red. The Hubble parameter (green), scaled by $10\times$ for readability. The co-moving Hubble radius (blue). Model parameters: $c=2$, $\alpha_{0}=\alpha=1$.
@@
Figure 2: Curvature invariants computed from (3.11). The Ricci scalar is plotted in green and the Kretchmann scalar is plotted in red. Model parameters: $c=2$, $\alpha_{0}=\alpha=1$.
@@
Figure 3: Energy conditions from (3). Plot of energy density $\rho$ (blue), $\rho+p$ (yellow), $|p|$ (red) and $\rho+3p$ (green). Model parameters: $c=2$, $\alpha_{0}=\alpha=1$.
All of the energy conditions are violated. NEC violation is identified by the dipping of the yellow curve ($\rho+p$) below the vertical axis, both here, and in all of the proceeding energy condition plots. The NEC asymptotically approaches saturation in the early and late universe since $\rho+p=-2\operatorname{sech}^{2}{t}\rightarrow 0$ as $t\rightarrow\pm\infty$. It is easy to see the ANEC (Eq. 1.1) is also violated in this scenario. We will see that the ANEC violation is a generic feature of flat eternal models. In the late universe, the SEC remains violated allowing continued acceleration, as indicated by the negativity of the green curve.
@@ IV.1 Transcendental bounce
The above scale factor describes a universe that contracts, bounces and then expands (similar to the familiar case of closed de Sitter).  [^10] A solution with this form was found in a non-local higher-derivative gravity in Biswas et al. (2006, 2012), which was argued to be free of pathologies, which if true would be quite remarkable given the level of energy condition violation (see Fig. 6). The Hubble parameter is
@@
$\int^{t}\frac{a\left(\zeta\right)d\zeta}{\sqrt{\left(a\left(\zeta\right)\right)^{2}+1}}=\alpha\operatorname{arctanh}\left(\frac{a_{0}\sinh\left(\frac{t}{\alpha}\right)}{\sqrt{1+a_{0}^{2}+a_{0}^{2}\sinh\left(\frac{t}{\alpha}\right)^{2}}}\right)\,$, (4.19)
@@
$\int^{t}a\left(\zeta\right)d\zeta=a_{0}\alpha\sinh\left(\frac{t}{\alpha}\right)\,.$ (4.20)
@@
Figure 4: Evolutionary behavior for Eq. 4.16. The scale factor amplified by an order of magnitude (red). The Hubble parameter amplified by an order of magnitude (green). The co-moving Hubble radius (blue). Parameters: $a_{0}=\alpha=1$.
@@
Figure 6: Energy conditions from Eq. IV.1. Plot of energy density $\rho$ (blue), $\rho+p$ (yellow), $|p|$ (red) and $\rho+3p$ (green). Model parameters: $a_{0}=\alpha=1$.
@@ IV.2 Polynomial bounce
Proposition 4.
Every smooth, non-constant scale factor $a>0$ of a geodesically complete GFRW spacetime must inflate for at least some period of time.
Before the proof is stated, a lemma from analysis:
Lemma 5.
Let $f:\mathbb{R}\rightarrow\mathbb{R}^{+}$ be twice differentiable with $f^{\prime\prime}\leq 0$ over all $\mathbb{R}$. Then $f$ must be a constant function: thus no globally concave positive non-constant real functions exist.
@@
First, assume that $f$ is not a constant function. There are now two cases to examine for $f^{\prime}$: for some $t_{0}\in\mathbb{R}$ either $f^{\prime}\left(t_{0}\right)>0$ or $f^{\prime}\left(t_{0}\right)<0$.
Let $f^{\prime}\left(t_{0}\right)>0$. By $f^{\prime\prime}\leq 0$ one reaps $f^{\prime}$ is non-increasing. Starting from the Fundamental Theorem of Calculus one can calculate
$f\left(t\right)=f\left(t_{0}\right)-\int^{t_{0}}_{t}f^{\prime}\left(s\right)ds<f\left(t_{0}\right)-f^{\prime}\left(t_{0}\right)\left(t_{0}-t\right)$ (4.24)
In the limit
$\lim_{t\rightarrow-\infty}f\left(t\right)<\lim_{t\rightarrow-\infty}f\left(t_{0}\right)-f^{\prime}\left(t_{0}\right)\left(t_{0}-t\right)=$-$\infty$ (4.25)
thus $f$ cannot have range $\mathbb{R}^{+}$.
Likewise assume $f^{\prime}\left(t_{0}\right)<0.$ In a similar vein
$f\left(t\right)=f\left(t_{0}\right)+\int^{t}_{t_{0}}f^{\prime}\left(s\right)ds<f\left(t_{0}\right)+f^{\prime}\left(t_{0}\right)\left(t-t_{0}\right)\,$ (4.26)
again taking
$\lim_{t\rightarrow\infty}f\left(t\right)<\lim_{t\rightarrow\infty}f\left(t_{0}\right)+f^{\prime}\left(t_{0}\right)\left(t-t_{0}\right)=$-$\infty$ (4.27)
Thus $f^{\prime}=0$ over all $\mathbb{R}$: hence $f$ is a constant function. ∎
Now the proof of Cor. 4 can be stated:
@@
Assume that $a^{\prime\prime}\leq 0$ over all $\mathbb{R}$; by Lemma 5 this implies $a$ is a constant function: a contradiction. Thus there is at least one point where $a^{\prime\prime}>0$; hence the spacetime is inflationary. ∎
Considering Corollary 4 and Defn. 1 we find non-trivial complete cosmologies must be inflationary Easson and Lesnefsky (2024). [^11] We pause here to acknowledge the celebrated work of Alexei Starobinsky who was motivated to find non-singular cosmological solutions by having an early-time accelerated expansion Starobinsky (1980), opening the window for the theory of the inflationary universe, and aligning with Cor.4.
IV.2.1 Polynomial model
Curvature and GR tensors can be computed in generality for a degree $n$ “appropriate” scale factor model per Defn. 3. To evaluate the aforementioned class of models said polynomial constants $\left\{a_{\ell}\right\}_{\ell=0}^{n}$ must be evaluated. The space of polynomial constants $\left\{a_{\ell}\right\}^{n}_{\ell=0}$ are not unconstrained–at minimum the requirements that the scale factor has no real roots yields a cursory set of limitations-namely that $a_{n}>0$ and $n$ even, proemially. It is well known that discriminants derived from radical methods can be utilized to reap polynomials with no real roots up to quartic degree, but more advanced methods including Galois theory, must be employed for quintic degree and higher - see Dummit and Foote (2004) for a standard graduate reference. The equations below are simple evaluations of well known curvature and GR tensors of GFRW models applied to an “appropriate” polynomial scale factor. All equations will be left in “summation form”, seeing as how there is no canonical way to expand said equations without further evaluation.
Recapitulating Eq. 4.23 of an “appropriate” polynomial
$\mathbb{R}\left[t\right]\ni p\left(t\right)=\sum_{\ell=0}^{n}a_{\ell}t^{\ell}$ (4.28)
in FRW [^12] A propos Thm. 1, the geodesic completeness of a GFRW is determined by its scale factor. For purposes of simplicity, a flat ($k=0$) FRW is utilized for computation. $\mathbb{R}^{1}_{1}\times_{p}\mathbb{R}^{n-1}$ where $\mathbb{R}\left[t\right]$ is the ring of polynomials over indeterminate $t$ taken to be the privileged time coordinate foliated in $\mathbb{R}^{1}_{1}$ and $a_{\ell}\in\mathbb{R}$ are the real coefficients of said polynomial.
Applying $p\left(t\right)$ to the formalism, one can calculate Kretchhmann and Ricci scalars to be
$K=$ $\frac{12\left(\left(\sum_{k=0}^{n}ka_{k}t^{k-1}\right)^{4}+\left(\sum_{k=0}^{n}(k-1)ka_{k}t^{k-2}\right)^{2}\left(\sum_{k=0}^{n}a_{k}t^{k}\right)^{2}\right)}{\left(\sum_{k=0}^{n}a_{k}t^{k}\right)^{4}}\,$, (4.29)
$R=$ $\frac{6\left(\left(\sum_{k=0}^{n}ka_{k}t^{k-1}\right)^{2}+\left(\sum_{k=0}^{n}(k-1)ka_{k}t^{k-2}\right)\sum_{k=0}^{n}a_{k}t^{k}\right)}{\left(\sum_{k=0}^{n}a_{k}t^{k}\right)^{2}}\,.$ (4.30)
The non-zero components of the Einstein tensor can be calculated
$G_{tt}=$ $\frac{3\left(\sum_{k=0}^{n}ka_{k}t^{k-1}\right)^{2}}{\left(\sum_{k=0}^{n}a_{k}t^{k}\right)^{2}}\,$, (4.31)
$G_{ii}=$ $-\left(\sum_{k=0}^{n}ka_{k}t^{k-1}\right)^{2}-2\left(\sum_{k=0}^{n}(k-1)ka_{k}t^{k-2}\right)\sum_{k=0}^{n}a_{k}t^{k}\,.$ (4.32)
The values of $\rho=-G^{t}_{t}$, $p=G^{i}_{i}$, and all subsequent energy conditions are trivially read from the previous equation. No plots can be generated until $\left\{a_{\ell}\right\}^{n}_{\ell=0}$ are specified.
For the sake of completeness, the Hubble parameter is given
$H=\frac{\sum_{k=0}^{n}ka_{k}t^{k-1}}{\sum_{k=0}^{n}a_{k}t^{k}}\,.$ (4.33)
IV.2.2 Quadratic model
@@
The absence of curvature singularities is shown from the curvature invariants plotted in Fig. 8. Additionally, the geodesics can be computed in closed form, and the existence of said solution for all time can be shown - see Appendix D.
@@
Figure 9: Energy conditions from Eq. IV.2.2. Plot of energy density $\rho$ (blue), $\rho+p$ (yellow), $|p|$ (red) and $\rho+3p$ (green). Model parameters: $a_{0}=b=1$, $c=2$.
@@
In this model, and others we consider, the NEC is violated during a finite time interval, $t_{-}<t<t_{+}$. One may break up the ANEC integral into the asymptotic regions $_{-}, t>t+t>t_{+}$ where other energy conditions ( e.g. NEC, WEC and DEC) are satisfied, and the NEC violating region. In this case the ANEC condition Eq. 1.1 may be recast as Giovannini (2017):
$\left|\int_{t_{-}}^{t_{+}}(\rho+p)d\lambda\right|\leq\int_{-\infty}^{-t_{-}}(\rho+p)d\lambda+\int_{t_{+}}^{+\infty}(\rho+p)d\lambda\,$, (4.40)
@@
In this case the ANEC integral gives:
$\int_{-\infty}^{\infty}(\rho+p)\,d\lambda$ $=\left(\frac{2a_{0}(2a_{0}t+b)}{(b^{2}-4a_{0}c)(a_{0}t^{2}+bt+c)}-\frac{8a_{0}^{2}\tan^{-1}\left(\frac{2a_{0}t+b}{\sqrt{4a_{0}c-b^{2}}}\right)}{(4a_{0}c-b^{2})^{3/2}}\right$.
$\quad\left.-\frac{2a_{0}t+b}{(a_{0}t^{2}+bt+c)^{2}}\right)\Big{|}_{-\infty}^{\infty}\leq 0\,$, (4.41)
which is negative for generic parameter values indication the ANEC is violated.
In the late universe, only the SEC is violated allowing continued acceleration, as indicated by the negativity of the green curve. This too is an eternally inflating, bouncing cosmology similar to the behavior of Eq. 4.16, but it is somewhat better-behaved with respect to energy conditions as the NEC violation may be restricted to a moment in the very early universe.  [^13] A modern example of such a bounce with momentary NEC violation is the S-brane model Brandenberger et al. (2014).
IV.2.3 Models from limiting curvature
One method for building such non-singular bounces stems from the limiting curvature hypothesis (LCH) which stipulates the existence of minimal length (or maximal curvature) scale in nature, presumably somewhere near the Planck length Markov (1982, 1987), although one may introduce new physics at lower energy scales using this method. By construction, curvature invariants are limited in an action resulting in non-singular solutions. The LCH has been used to successfully construct many non-singular cosmological solutions Mukhanov and Brandenberger (1992); Brandenberger et al. (1993); Moessner and Trodden (1995); Brandenberger (1995); Brandenberger et al. (1998); Easson and Brandenberger (1999); Easson (2003a, 2007); Chamseddine and Mukhanov (2017a) in a variety of settings, as well as non-singular black hole solutions Frolov et al. (1989, 1990); Morgan (1991); Trodden et al. (1993); Easson (2003b, 2018); Chamseddine and Mukhanov (2017b); Brandenberger et al. (2021); Frolov and Zelnikov (2022).
@@ IV.2.3 Models from limiting curvature
for constants $c>0,b>0$, which is capable of smoothly connecting two matter dominated phases having $a\propto t^{2/3}$ via a NEC violating bounce at $t=0$ where $\ddot{a}>0$. The schematic behavior of the cosmological quantities and the energy conditions are very similar to the polynomial model analyzed above; however, in the present case, even the SEC is obeyed in both the early and late (non-accelerating) universe. The ANEC integral Eqs. 1.1 and 4.40 may be calculated in terms of elliptic functions:
$\int_{-\infty}^{\infty}(\rho+p)\,d\lambda$ $=\frac{2a_{0}(2a_{0}t+b)}{(b^{2}-4a_{0}c)(a_{0}t^{2}+bt+c)}-\frac{8a_{0}^{2}\tan^{-1}\left(\frac{2a_{0}t+b}{\sqrt{4a_{0}c-b^{2}}}\right)}{(4a_{0}c-b^{2})^{3/2}}$
$\quad-\frac{2a_{0}t+b}{(a_{0}t^{2}+bt+c)^{2}}\Big{|}_{-\infty}^{\infty}\leq 0\,$, (4.43)
which is negative for generic parameter values.